On January 1, 2018 Friendly Farm Company purchased a new machine at a cost of $350,000. The machine has an estimated useful life of 4 years or 100,000 hours and residual value of $30,000. The machine will be used 30,000 hours in year 1, 40,000 hours in year 2, 20,000 hours in year 3 and 10,000 hours in year 4.

Requirements:

IN EXCEL FORMAT- Prepare a depreciation schedule using each of the three methods: Straight line, Units of Production and Double Declining Balance.
Answer #1

Answers

  • Straight Line method

A

Cost

$          350,000.00

B

Residual Value

$            30,000.00

C=A - B

Depreciable base

$          320,000.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$            80,000.00

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$           350,000.00

$            80,000.00

$       270,000.00

$             80,000.00

2

$           270,000.00

$            80,000.00

$       190,000.00

$           160,000.00

3

$           190,000.00

$            80,000.00

$       110,000.00

$           240,000.00

4

$           110,000.00

$            80,000.00

$         30,000.00

$           320,000.00

  • Units of production

A

Cost

$          350,000.00

B

Residual Value

$            30,000.00

C=A - B

Depreciable base

$          320,000.00

D

Usage

                   100,000

E = C/D

Depreciation per …..

$                    3.200

Year

Book Value

Usage

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$           350,000.00

                     30,000

$         96,000.00

$           254,000.00

$        96,000.00

2

$           254,000.00

                     40,000

$       128,000.00

$           126,000.00

$      224,000.00

3

$           126,000.00

                     20,000

$         64,000.00

$             62,000.00

$      288,000.00

4

$              62,000.00

                     10,000

$         32,000.00

$             30,000.00

$      320,000.00

  • Double Declining balance

A

Cost

$          350,000.00

B

Residual Value

$            30,000.00

C=A - B

Depreciable base

$          320,000.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$            80,000.00

F=E/C

SLM Rate

25.00%

G=F x 2

DDB Rate

50.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$           350,000.00

50.00%

$       175,000.00

$           175,000.00

$      175,000.00

2

$           175,000.00

50.00%

$         87,500.00

$             87,500.00

$      262,500.00

3

$              87,500.00

50.00%

$         43,750.00

$             43,750.00

$      306,250.00

4

$              43,750.00

$         13,750.00

$             30,000.00

$      320,000.00